# Careful finite element simulations of cold rolling with accurate through-thickness resolution and prediction of residual stress

### v1

@@ Abstract
In this paper, we carefully develop a finite element (FE) model that gives accurate through-thickness predictions of stress and strain distributions during the steady-state cold rolling process. These through-thickness predictions unveil an oscillatory pattern that is shown to have important consequences for residual stress in the rolled sheet. We believe this is the first time that through-thickness FE results have been accurately validated by comparison to non-FE results, in this case by comparison to a recent analytical model of through-thickness variation in cold rolling. While we use here the ABAQUS commercially available FE software, our observations are relevant to all FE simulations of cold rolling. Care is taken by considering both convergence in number of elements through thickness, convergence to a steady state, and the avoidance of other numerical artefacts such as shear locking and hourglassing. We find that previous FE models of cold rolling are usually woefully under-resolved through-thickness; e.g. using 10 elements through-thickness, while we require 60 here for convergence. Convergence of roll force and roll torque, used in previous studies to validate models, are shown to be poor indicators of through-thickness convergence. We also show that the through-thickness oscillatory pattern may have important consequences for predicting curvature during asymmetric rolling.
@@ 1 Introduction
Energy used in industrial activity produces 40% of global $\mathrm{CO}_{2}$ emissions, the majority of which are caused by the manufacture of primary materials such as steel and aluminium [1]. Steel manufacturing is particularly problematic when one considers that crude steel production grew by approximately 900% between 1950 and 2023 [2], and that global steel demand is expected to grow by another 2.9% by 2025 [3]. The underlying mechanics of metal forming processes are not yet fully understood, and hence the problem of efficiently modifying the tool path based on the measured error is still unsolved [4]. As a result of these misunderstandings, liquid metal must be overproduced to ensure final product quality. For manufacturers this increases costs and reduces productivity [5] but, as liquid metal production is the most energy and $\mathrm{CO}_{2}$-intensive processing step, these knowledge gaps also have significant implications for sustainability. Reducing liquid metal demand through process improvements could have a significant impact on energy use and $\mathrm{CO}_{2}$ emissions [6]. In this work we will focus on cold rolling, a common steel manufacturing method.
@@
Figure 1: Schematic of the metal rolling process. The sheet moves from left to right and its thickness is reduced by the roller. Symmetry about the centre of the sheet is assumed.
The most common method of assessing cold rolling is through the application of mathematical models, which can be used to optimise system design and tighten process specification [1, 7]. Most traditional slab-type mathematical models are based on the classical Orowan [8] approach. In general, slab models assume that the stresses are vertically homogeneous through the thickness of the sheet, that each vertical slab of material remains vertical and that the sheet does not stick to the rollers throughout the rolling process. While this very simple mathematical description can yield reasonably accurate predictions of the roll force and torque required for a desired thickness reduction, they disregard through-thickness effects [9, 10]. These through-thickness distributions are significant for downstream properties such as residual stress which in turn informs the overall quality of the final product. It is therefore desirable to enhance the understanding of the through-thickness stress and strain in the sheet.
Through-thickness variations in stress and strain can be attained through more advanced modelling techniques such as finite element (FE) analysis. The FE method can give high-resolution predictions of stresses and strains in the roll gap. In contrast to existing analytical models, FE models also predict stress distributions outside the roll gap. Recent FE simulations of cold rolling by the current authors [11] show oscillatory stress and strain patterns. These distributions appear to have important consequences for cold-rolling residual stresses [12, 13, 11], while slab models fail to capture these variations. Of course, the FE method is still an approximate solution technique with its own limitations, such as slow computational time. This makes the FE method unsuitable for real-time process control [14]. Convergence issues for extreme parameter values also limit the range of possible explorations that can be considered during FE simulations. For example, elasticity is often neglected by assuming an infinite Young’s modulus, $E$, in analytical modelling for simplification, but this is impossible to reproduce in certain elastic-plastic FE softwares such as ABAQUS. Analytical approaches such as asymptotics-based models have the potential to be used in online control, assuming they can output sufficiently accurate results. Although not directly useful for online control, FE outputs can be used as a benchmark for these quicker-to-compute analytical models to ensure that the analytical modelling gives satisfactory outputs. The FE method allows for a range of parameter regimes to be explored at a relatively low cost compared to experimental results [10]. To act as a benchmark for through-thickness predictions, the FE model used needs a sufficiently fine mesh, among other model settings, which we will see later in this work. The overarching aim of this paper is to provide a reliable FE model that accurately captures the through-thickness stress and strain variations arising in the sheet during rolling, which can guide the development of simpler, faster mathematical models. As we will see, such a reliable FE model is absent from existing literature.
@@
In Section 2, we provide a brief overview of existing analytical and FE cold-rolling models that consider through-thickness stress and strain variations, while an in-depth description of the ABAQUS/Standard and ABAQUS/Explicit FE formulations may be found in the supplementary material. In Section 3 we provide details of the implicit and explicit FE cold rolling models. In Section 4 we conduct a mesh study and compare implicit and explicit in terms of their ability to accurately predict through-thickness stress variations. We show that roll force and torque are poor determinants of the optimal mesh size and of the steady state condition. In Section 5, we compare the simulated FE results with a new analytical model from Erfanian et al. [13]. We also show that the residual stress in the sheet is related to the mechanics inside the roll gap. In Section 6 we draw conclusions and briefly discuss the future plans for our FE simulations.
@@ 2 Literature review
A range of commerical FE packages are available for modelling metal deformation, such as Ansys, DEFORM, LS-Dyna, MSC Marc and ABAQUS. FE simulations reported in this work were conducted using ABAQUS 2021 [15]. Within the ABAQUS package there are two numerical solution types: implicit and explicit. For the cold rolling process it is not immediately clear which solution type is optimal as each type offers different advantages and disadvantages. The explicit solver is often the preferred choice for modelling metal forming processes due to its computational speed and robust handling of complex contact conditions [16, 17, 18, 19]. A key disadvantage of choosing the explicit solver is the absence of equilibrium convergence enforcement which can lead to error accumulation [20]. The potential for explicit solvers to accumulate error implies that implicit solvers provide stronger results [21]. Implicit solvers however typically consume more computational resources and can struggle to handle complex contact [19]. In Section 3 we will assess both implicit and explicit solvers for modelling cold rolling, with the specific aim of accurately capturing through-thickness variations of stress and strain quantities.
@@
Asymptotic analysis is useful for modelling physical processes where the smallness (or largeness) of a fundamental process parameter can be exploited [22]. This type of analysis is particularly useful for cold rolling, where the roll-gap aspect ratio is typically large [14], i.e., $\varepsilon=h_{0}/L\ll 1$, where $h_{0}$ is the initial half-thickness of the sheet and $L\approx\sqrt{2R\Delta h}$ is the horizontal length of the roll gap (see Figure 1). Domanti and McElwain [23] assume a small reduction in the sheet’s thickness (i.e., $r=\Delta h/h_{0}\ll 1$) along with the large-aspect-ratio assumption. These assumptions enable an asymptotic analysis to be carried out to obtain accurate solutions in the limits as $\varepsilon\rightarrow 0$ and $r\rightarrow 0$. However, these assumptions limit the range of validity of the analytical model. Erfanian et al. [13] relax the assumption that the reduction of the sheet thickness is small and also resolve behaviour over two distinct length scales associated with $h_{0}$ and $L$. By contrast, Domanti and McElwain [23] and other authors [see, for example, 14] consider variations on a length scale of $L$ only. The multiple-scales analysis of Erfanian et al. [13] recovers slab theory at leading order, and at higher orders predicts through-thickness oscillations in stress and velocity in the roll gap. These through-thickness variations have also been captured by the current authors in recent FE simulations [11]. The Erfanian et al. [13] model is limited to rigid-perfectly-plastic materials, thus elastic and strain-hardening effects are neglected. FE models can integrate elastic-plastic material behaviour which captures elasticity and strain hardening. Comparing FE outputs to analytical models such as the Erfanian et al. [13] model can give an understanding of what mathematical assumptions are valid or invalid.
[24] 2D 10 RP Mesh distortion, relative velocities, normal/shear strain rates, PEEQ
[25] 2D 12 Six-node triangular EVP Effective stress
[26] 2D 10 Isoparametric quad full integration RP Mesh distortion, PEEQ rate, normal stress
[27] 3D 8–10 8-node brick EP Velocity flow patterns
[28] 2D 8 Quad, reduced-integration EP Stress components
[29] 3D 6 EP Mesh distortion
[30] 2D 14 Quad RP PEEQ rate, pressure
[31] 3D 3 8-node brick EP Incremental strain components
[32] 2D 20 Quad EP Mesh distortion, PEEQ
[33] 2D 8 RP PEEQ, PEEQ rate Asymmetric
[34] 2D 18 4-node EoP Stress (centre & surface only)
[35] 2D 18 Quad EVP Mesh distortion, PEEQ, horizontal stress Asymmetric
[36] 3D 10 8-node brick, reduced-integration EP Effective stress, PEEQ
[37] 2D 8 Rectangular RP PEEQ
[38] 2D 16–18 Plane-strain 4-node VP Temperature, strain, strain rate, effective stress
[9] 2D Isoparametric quad elements EP PEEQ
[39] 3D 10 8-node brick reduced-integration EP Vertical, shear stress components
[40] 3D 12 8-node brick reduced-integration temperature-coupled EP Normal stress, pressure, displacements (edge only)
[41] 2D 16 Square isoparametric RP PEEQ, displacements, velocities
[42] 2D 36 3-noded triangular RP PEEQ, temperature
[43] Both 2D: 18-30, 3D: 11 2D: Plane-strain 4-node reduced-integration, 3D: 8-node reduced-integration EP Stress, velocity components Symmetric and asymmetric
[44] 2D 20 4-node isoparametric EP Temperature, heat flux
[45] 2D 16 4-node square elements EP Longitudinal, shear and effective strain, longitudinal and von Mises stress
[11] 2D 12–100 Plane-strain 4-node reduced-integration EP Vertical velocity, von Mises
@@
Table 1 shows the common through-thickness variations from existing FE models. As shown in Table 1, comparing through-thickness variations from various works is difficult given the distinct differences in modelling criteria and assumptions (e.g., 2D or 3D simulation, number of elements through-thickness ($2N_{e}$), element type, etc.) made by authors. The implication of this lack of comparability means that there is no clear mechanism for novices to recognise what constitutes a “good” FE rolling simulation. While several authors report shear stress along the roller-sheet interface [16, 46, 14], Table 1 shows that few [28, 39, 43] provide the through-thickness shear stress variation. Given the potential implications of oscillating shear stress on downstream quantities such as residual stress [12, 11], a “good” FE model should be capable of accurately predicting the oscillatory shear stress (and other quantities) through the thickness of the sheet.
@@ 3.1 Model geometry
Since the metal sheet’s width is much larger than its thickness in cold rolling in general, we consider an idealised two-dimensional geometry by assuming plane-strain conditions apply [46]. Readers are referred to Figure 1 for a diagram of the model geometry. We consider a rigid roller of radius $R=257.45\,$\mathrm{m}\mathrm{m}$$, rolling a sheet of initial half-thickness $h_{0}=2\,$\mathrm{m}\mathrm{m}$$, and impose an approximate half-thickness reduction of $\Delta h=0.5\,$\mathrm{m}\mathrm{m}$$. This corresponds to a roll-gap aspect ratio of $\varepsilon=h_{0}/L=0.125$, where $L\approx 16.05\,$\mathrm{m}\mathrm{m}$$ is the approximate horizontal length of the roll gap.
@@ 3.2 Material properties
We consider a sheet of elastic-plastic mild steel (grade DC04), a common material used in cold metal forming [49]. The DC04 material deforms elastically up to an initial yield stress of $Y=477.2\,$\mathrm{M}\mathrm{P}\mathrm{a}$$, and the yield stress increases to $650.25\,$\mathrm{M}\mathrm{P}\mathrm{a}$$, at a true plastic strain of 1.1, as a result of strain hardening. The Young’s modulus and Poisson’s ratio are taken to be $E=206.3\,$\mathrm{G}\mathrm{P}\mathrm{a}$$ and 0.3 respectively. For all explicit analyses, the density of the material is defined as $\rho=7831.3\,$\mathrm{k}\mathrm{g}\,\mathrm{m}^{-3}$$.
@@ 3.3 General FE model conditions
The sheet’s horizontal centre line (i.e., the symmetry axis in Figure 1) is constrained vertically to enforce the expected symmetry. This modelling choice reduces the computational intensity of the problem, and in turn facilitates the use of a finer mesh. The centre of the roller is defined as a reference point and boundary conditions, depicting the movement of the roller, are applied to this location. Initially, the roller and the sheet are not in contact, but are set up such that the roller is positioned above the sheet. In the bite step the roller is slowly translated vertically to indent the sheet, but no rotation of the roller occurs in the bite step. A bite step is used here, instead of starting the strip in front of the roll gap and feeding it in in some way, as the bite-then-roll simulations showed fewer initial contact issues and converged more quickly to steady state (as also described by Minton [43]). The bite step is performed via a velocity boundary condition at the roller’s reference point. In the rolling step, the roller reference point is fixed so its only degree of freedom is rotation in the rolling direction, and the roller is given a tangential velocity of $R\Omega=1287.25\,$\mathrm{m}\mathrm{m}\,\mathrm{s}^{-1}$$. Note that the above boundary conditions are used for all analyses below unless otherwise stated.
As the formulation differs for implicit and explicit simulations the model specifications understandably diverge. Table 3 provides model settings unique to implicit and explicit simulations.
@@
The amplitude curve information given in Table 3 relates to the application of loading on the roller during the rolling step. Implicit simulations detailed here apply an instantaneous $\Omega=5~{}$\mathrm{r}\mathrm{a}\mathrm{d}\,\mathrm{s}^{-1}$$ while explicit simulations smoothly ramp loading to $\Omega=5~{}$\mathrm{r}\mathrm{a}\mathrm{d}\,\mathrm{s}^{-1}$$ over a fraction of the rolling step.
In our analyses we compare results from the four-node, plane-strain elements. As detailed in Table 3, implicit analyses include comparisons of full integration (CPE4), reduced integration (CPE4R), and incompatible (CPE4I) type elements. Since full integration (CPE4) and incompatible type (CPE4I) elements are not offered by ABAQUS/Explicit only reduced integration (CPE4R) elements are considered. Meshing is applied using a global seed formulation with a structured meshing technique. Varying levels of mesh density are considered and the optimal mesh is discussed in detail in Section 4.

### v2

@@ Abstract
Metal rolling is a widespread and well-studied process, and many finite-element (FE) rolling simulations can be found in the scientific literature. However, these FE simulations are typically limited in their resolution of through-thickness variations. In this paper, we carefully assess the accuracy of a number of FE approaches, and find that at least 60 elements through-thickness are needed to properly resolve through-thickness variation; this is significantly more than is used elsewhere in the metal rolling literature. In doing so, we reveal an oscillatory stress pattern, which is not usually observed in simulations but which we can validate by comparison with recent analytical work, and which is completely deterministic, not arising from numerical noise or error. We show that these oscillations contribute to the formation of residual stress and may help predict curvature in asymmetric rolled sheets, a phenomenon which is currently not well understood. Accurate through-thickness variation of stress and strain would also have implications for modelling microstructure evolution, damage, and surface finish.
@@ 1 Introduction
Figure 1: Schematic of the metal rolling process. The sheet moves from left to right and its thickness is reduced by the roller. Symmetry about the centre of the sheet ($z=0$) is assumed, so only the top roller and top half of the sheet are modelled in this work.
Mathematical models of the cold-rolling process have been used for many years to make predictions, optimise system design and tighten process specification [1, 2, 3]. Waste reduction is also an important motivator in this context as the production of steel and aluminium grows year-on-year and is a major contributor to global energy consumption and CO2 emissions [4, 5, 6].
In this paper we re-visit the task of modelling cold rolling using finite element (FE) analysis, with a particular focus on through-thickness variations. We show that rolled sheets contain a rich and detailed stress pattern that varies repeatedly across the sheet thickness and that has not been properly resolved previously. These oscillations do not arise from noise, vibrations or uncertainty, but rather are completely deterministic. From a modelling point of view, we assess the accuracy of different FE approaches for modelling these through-thickness variations and recommend an approach that can fully resolve stress patterns. From a physics and process engineering point of view, we fully resolve and validate these patterns and outline their implications for downstream effects like residual stress and curvature.
Early mathematical descriptions of cold rolling used a “slab” approach pioneered by Orowan [1]. Slab models assume that the sheet can be thought of as a series of vertical slabs which remain vertical throughout the rolling process, and that stresses are uniform across the sheet thickness. While this simple mathematical description can yield reasonably accurate predictions of the roll force and torque required for a given thickness reduction, they automatically disregard any through-thickness effects [7]. Experimental observations of deformation led to pictures like the one shown in Figure 2(a), with zones of thinning, thickening and restricted deformation. Through-thickness variations in stress and strain can in principle be described with more advanced modelling techniques such as FE analysis [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. However, we show that although the mesh resolution typically used in the scientific literature is high enough for most purposes, it is not high enough to properly resolve the oscillatory stress pattern of interest here. Using the highly careful FE model developed here, we will show below that the deforming and non-deforming zones inside the roll gap can exhibit a much more detailed pattern than previously reported (see, for example, Figure 2(b) as opposed to Figure 2(a)).
Refer to captionIIIIIIIIV
Figure 2: Deformation zones in a metal sheet during rolling: (a) deformation zones reported by [19] using data from [20]. Region I has “restricted deformation”, II has thinning, and III and IV have thickening. (b) PEEQ rate predicted here. Grey zones are not flowing, even if they are at yield, while darker zones are shearing.
@@
In Section 2, we provide a brief overview of existing analytical and FE cold-rolling models that consider through-thickness stress and strain variations, while an in-depth description of the ABAQUS/Standard and ABAQUS/Explicit FE formulations may be found in A. In Section 3 we provide details of the implicit and explicit FE cold rolling models. In Section 4 we conduct a mesh study and compare implicit and explicit in terms of their ability to accurately predict through-thickness stress variations. We show that roll force and torque are poor determinants of the optimal mesh size and of the steady state condition. In Section 5 we present new physical phenomena predicted by our FE simulations, and show that the residual stress in the sheet is related to oscillatory mechanics inside the roll gap. In Section 6 we provide further validation by comparison with a new analytical model from Erfanian et al. [21]. In Section 7 we draw conclusions and briefly discuss the future plans for our FE simulations.
@@ 2 Literature review
FE simulations reported in this work were conducted using ABAQUS 2021 [22]. Within the ABAQUS package there are two numerical solution types: implicit and explicit. For the cold rolling process it is not immediately clear which solution type is optimal as each type offers different advantages and disadvantages. The explicit solver is often the preferred choice for modelling metal forming processes due to its computational speed and robust handling of complex contact conditions [23, 24, 25, 26]. A key disadvantage of choosing the explicit solver is the absence of equilibrium convergence enforcement which can lead to error accumulation [27]. The potential for explicit solvers to accumulate error implies that implicit solvers provide stronger results, since convergence is ensured at each time increment in implicit analyses [28]. However, implicit solvers typically consume more computational resources and can struggle to handle complex contact [26]. In Section 3 we will assess both implicit and explicit solvers for modelling cold rolling, with the specific aim of accurately capturing through-thickness variations of stress and strain quantities.
@@
Asymptotic analysis is useful for modelling physical processes where the smallness (or largeness) of a fundamental process parameter can be exploited [29]. This type of analysis is particularly useful for cold rolling, where the roll-gap aspect ratio is typically large [30], i.e., $1/\varepsilon=L/h_{0}\gg 1$, where $h_{0}$ is the initial half-thickness of the sheet and $L=\sqrt{2R\Delta h}$ is the approximate horizontal length of the roll gap (see Figure 1). Domanti and McElwain [31] assume a small reduction in the sheet’s thickness (i.e., $r=\Delta h/h_{0}\ll 1$) along with the large-aspect-ratio assumption. These assumptions enable an asymptotic analysis to be carried out to obtain accurate solutions in the limits as $\varepsilon\rightarrow 0$ and $r\rightarrow 0$. However, these assumptions limit the range of validity of the analytical model. Erfanian et al. [21] relax the assumption that the reduction of the sheet thickness is small and also resolve behaviour over two distinct length scales associated with $h_{0}$ and $L$. By contrast, Domanti and McElwain [31] and other authors [see, for example, 30] consider variations on a length scale of $L$ only. The multiple-scales analysis of Erfanian et al. [21] recovers slab theory at leading order, and at higher orders predicts through-thickness oscillations in stress and velocity in the roll gap. These through-thickness variations have also been captured by the current authors in recent FE simulations [32]. The Erfanian et al. [21] model is limited to rigid-perfectly-plastic materials, thus elastic and strain-hardening effects are neglected. FE models can integrate elastic-plastic material behaviour which captures elasticity and strain hardening. Comparing FE outputs to analytical models such as the Erfanian et al. [21] model can give an understanding of what mathematical assumptions are valid or invalid.
@@
Table 1 shows the common through-thickness variations from existing FE models. As shown in Table 1, comparing through-thickness variations from various works is difficult given the distinct differences in modelling criteria and assumptions (e.g., 2D or 3D simulation, number of elements through-thickness ($2N_{e}$), element type, etc.) made by authors. The implication of this lack of comparability means that there is no clear mechanism for novices to recognise what constitutes a “good” FE rolling simulation. While several authors report shear stress along the roller-sheet interface [23, 33, 30], Table 1 shows that few [10, 17, 18] provide the through-thickness shear stress variation. Given the potential implications of oscillating shear stress on downstream quantities such as residual stress [34, 32], we argue that a “good” FE model should be capable of accurately predicting the oscillatory shear stress (and other quantities) through the thickness of the sheet.
@@
Li and Kobayashi [40] 2D 10 RP Mesh distortion, relative velocities, normal/shear strain rates, PEEQ
Thompson [8] 2D 12 Six-node triangular EVP Effective stress
Mori et al. [9] 2D 10 Isoparametric quad full integration RP Mesh distortion, PEEQ rate, normal stress
Liu et al. [43] 3D 8–10 8-node brick EP Velocity flow patterns
Yarita et al. [10] 2D 8 Quad, reduced-integration EP Stress components
Liu et al. [38] 3D 6 EP Mesh distortion
Mori and Osakada [44] 2D 14 Quad RP PEEQ rate, pressure
Liu et al. [41] 3D 3 8-node brick EP Incremental strain components
Lau et al. [11] 2D 20 Quad EP Mesh distortion, PEEQ
Yoshii et al. [12] 2D 8 RP PEEQ, PEEQ rate Asymmetric
Malinowski and Lenard [45] 2D 18 4-node EoP Stress (centre & surface only)
Richelsen [13] 2D 18 Quad EVP Mesh distortion, PEEQ, horizontal stress Asymmetric
Ghosh et al. [14] 3D 10 8-node brick, reduced-integration EP Effective stress, PEEQ
Gudur and Dixit [15] 2D 8 Rectangular RP PEEQ
Shahani et al. [46] 2D 16–18 Plane-strain 4-node VP Temperature, strain, strain rate, effective stress
Lenard [16] 2D Isoparametric quad elements EP PEEQ
Cawthorn et al. [17] 3D 10 8-node brick reduced-integration EP Vertical, shear stress components
Zhou et al. [47] 3D 12 8-node brick reduced-integration temperature-coupled EP Normal stress, pressure, displacements (edge only)
Misovic et al. [42] 2D 16 Square isoparametric RP PEEQ, displacements, velocities
Koohbor and Moaven [48] 2D 36 3-noded triangular RP PEEQ, temperature
Minton [18] Both 2D: 18-30, 3D: 11 2D: Plane-strain 4-node reduced-integration, 3D: 8-node reduced-integration EP Stress, velocity components Symmetric and asymmetric
Olaogun et al. [35] 2D 20 4-node isoparametric EP Temperature, heat flux
Tadić et al. [39] 2D 16 4-node square elements EP Longitudinal, shear and effective strain, longitudinal and von Mises stress
Flanagan et al. [32] 2D 12–100 Plane-strain 4-node reduced-integration EP Vertical velocity, von Mises
@@ 3.1 Model geometry
Since the metal sheet’s width–thickness ratio is, in general, large in cold rolling (for example, Hacquin et al. [49], Yao et al. [50, 51] and Li et al. [52] all employ a width–thickness ratio that is at least 100), we consider an idealised two-dimensional geometry by assuming plane-strain conditions apply [33, 53, 16]. Readers are referred to Figure 1 for a diagram of the model geometry. Only the top roller and the top half of the sheet are modelled in this work due to the expected symmetry about the symmetry axis at $z=0$. We consider a rigid roller of radius $R=257.45\,$\mathrm{m}\mathrm{m}$$, rolling a sheet of initial half-thickness $h_{0}=2\,$\mathrm{m}\mathrm{m}$$, and impose an approximate half-thickness reduction of $\Delta h=0.5\,$\mathrm{m}\mathrm{m}$$. This corresponds to $\varepsilon=h_{0}/L=0.125$, where $L=16.05\,$\mathrm{m}\mathrm{m}$$ is the approximate horizontal length of the roll gap.
@@ 3.2 Material properties
We consider a sheet of elastic-plastic mild steel (grade DC04), a common material used in cold metal forming [54]. The DC04 material deforms elastically up to an initial yield stress of $Y=477.2\,$\mathrm{M}\mathrm{P}\mathrm{a}$$, and the yield stress increases to $650.25\,$\mathrm{M}\mathrm{P}\mathrm{a}$$, at a true plastic strain of 1.1, as a result of strain hardening. The Young’s modulus and Poisson’s ratio are taken to be $E=206.3\,$\mathrm{G}\mathrm{P}\mathrm{a}$$ and $\nu=0.3$ respectively. For all explicit analyses, the density of the material is defined as $\rho=7831.3\,$\mathrm{k}\mathrm{g}\,\mathrm{m}^{-3}$$.
@@ 3.3 General FE model conditions
The sheet’s horizontal centre line (i.e., the symmetry axis in Figure 1) is constrained vertically to enforce the expected symmetry. This modelling choice reduces the computational intensity of the problem, and in turn facilitates the use of a finer mesh. The centre of the roller is defined as a reference point and boundary conditions, depicting the movement of the roller, are applied to this location. Initially, the roller and the sheet are not in contact, but are set up such that the roller is positioned above the sheet. In the bite step the roller is slowly translated vertically (via a velocity boundary condition) to indent the sheet, but no rotation of the roller occurs in the bite step. In the rolling step, the roller reference point is fixed so its only degree of freedom is rotation in the rolling direction, and the roller is given a tangential velocity of $R\Omega=1287.25\,$\mathrm{m}\mathrm{m}\,\mathrm{s}^{-1}$$. The bite step is used here, instead of starting the strip in front of the roll gap and feeding it in in some way, as the bite-then-roll simulations showed fewer initial contact issues and converged more quickly to steady state (as also described by Minton [18]). Note that the above boundary conditions are used for all analyses below unless otherwise stated.
As the formulation differs for implicit and explicit simulations the model specifications understandably diverge. Table 3 provides model settings unique to implicit and explicit simulations. The amplitude curve information given in Table 3 relates to the application of loading on the roller during the rolling step. Implicit simulations detailed here apply an instantaneous $\Omega=5~{}$\mathrm{r}\mathrm{a}\mathrm{d}\,\mathrm{s}^{-1}$$ while explicit simulations smoothly ramp loading to $\Omega=5~{}$\mathrm{r}\mathrm{a}\mathrm{d}\,\mathrm{s}^{-1}$$ over a fraction of the rolling step.
In the analysis below we compare results from the four-node, plane-strain elements. As detailed in Table 3, implicit analyses include comparisons of full-integration (CPE4), incompatible (CPE4I) and reduced-integration (CPE4R) element types. Since full-integration (CPE4) and incompatible type (CPE4I) elements are not offered by ABAQUS/Explicit only reduced-integration (CPE4R) elements are considered. Meshing is applied using a global seed formulation with a structured meshing technique. Varying levels of mesh density are considered and the optimal mesh is discussed in detail in Section 4.